Operations Report – February 11th

SOL: #7
Name of person filing report: Rajvi Patel
Non-nominal systems: None
Notes on non-nominal systems: Lower Hab Heater
The heater turned off again last night showing 4 orange flashes (1 short 3 long). I fixed it using the same method as earlier and informed Sergii about this. The heater has been working fine since then.
Spirit rover used: Not used
Opportunity rover used: EVA #8
Hours: 161.1
Beginning charge: 100%
Ending charge: 52%
Currently charging: No because of the late afternoon
Curiosity rover used: EVA #8
Hours: 266.1
Beginning charge: 100%
Ending charge: 46%
Currently charging: No because of the late afternoon
Perseverance rover used: Not used
General notes on rovers:
Summary of Hab operations: put notes here
WATER USE: 30.36 gallons
Water (static tank): 310.72 gallons
Static tank pipe heater (on or off): On
Static tank heater (On or off): On
Toilet tank emptied: Yes
Summary of internet: Nominal
Summary of suits and radios: Nominal. Suits and radios worked fine for the EVA #008.
-Regarding Suit#6: I have ran four set of tests.
1) Case a: 20 min run at max fans
0 min 12.71 V
20 min 12:18 V
40 min 6.92 V
60 min 4.13 V
80 min 3.16V

2) Case b: 3 min run at max fans
0 min 13.89 V
3 min 12:97 V
6 min 10.57 V
9 min 6.04 V
12 min 5.12 V

The fans started to slow down super quick after hitting 10 V and the rate of voltage reduction was quite fast from 10V to 5.3V. This was followed by a reduction in the rate of voltage decrease (e.g. approx. 0.5 V per second)
3) Case c: 20 min run at min fans
0 min 13.13 V
20 min 8.95 V
40 min 8.55 V
60 min 7.73 V
I think the battery needs to be replaced as it cannot last long while the fans are at max speed. Need your advice.
Summary of GreenHab operations: put notes here
WATER USE: 7.2 gallons
Heater: On
Supplemental light: 22:00 – 02:00
Harvest: None
Summary of ScienceDome operations: In the morning Daniel successfully carried out a PCR and gel electrophoresis to identify the microbes found in the DNA from the soil reported yesterday, results are yet to be analyzed. In the afternoon another PCR targeting biogeochemical cycling genes was carried out on the same DNA, currently waiting for the gel to finish running. DNA extractions took place concurrently on soil gathered in today’s EVA.
EVA.Dual split: Heat On
Summary of RAM operations: (Tools used, work done) : RAM was accessed to receive the medical supplies delivered by Sergii.
Summary of any observatory issues: None
Summary of health and safety issues: None
Questions, concerns, and requests to Mission Support: put notes here
The Toilet was emptied at midnight as the bad smell reached the Upper Hab and it was difficult for the crew to breathe. The sensor was still at ½ .

Copyright © The Mars Society. All rights reserved. | Main Site